# Fast LISA likelihood approximations by downsampling

### v1

@@ 4.3.1 Induced distribution noise
    Note that downsampling noise vanishes when $N_{\mathrm{s}}=N_{\mathrm{f}}$. Assume that the behaviour of the noise is such that as $N_{\mathrm{s}}$ is decreased, the noise increases or remains the same. Thus choosing some ‘relatively large’ $N_{\mathrm{s}}=\hat{N}_{\mathrm{s}}$ induces low posterior noise. This behaviour, and whether the ‘relatively large’ $\hat{N}_{\mathrm{s}}$ is an appropriate choice, are confirmed experimentally in subsequent steps.
@@
    Choose some (‘relatively large’) $\hat{N}_{\mathrm{s}}$, and compute both sampler noise & downsampling(+sampler) noise. That is, take some particular $\mathbf{k^{\prime}}$ (with $\hat{N}_{\mathrm{s}}$ 1’s) to define a posterior $p_{\mathbf{k^{\prime}}}$, produce the set of $K$ estimates $\{p^{i}_{\mathrm{smp}}\}$ using the sampler. The sampler noise is $\mu[X_{\mathrm{smp}}]$, where $X_{\mathrm{smp}}$ is given in equation (37). Then produce another set of $K$ estimates of posteriors $p_{\mathrm{DS+smp}}^{i}$ using different random realisations of $\mathbf{k}$ with $\hat{N}_{\mathrm{s}}$ 1’s. The downsampling (with sampler) noise is given by $\mu[X_{\mathrm{DS+smp}}]$.
@@
    Assuming an accurate and unbiased sampler, we have that $p^{i}_{\mathrm{smp}}\approx\mathbb{E}[p^{i}_{\mathrm{smp}}]=p_{\mathbf{k^{\prime}}}$. Thus if both $\mu[X_{\mathrm{DS+smp}}]\approx\mu[X_{\mathrm{smp}}]$ and $\sigma[X_{\mathrm{DS+smp}}]\approx\sigma[X_{\mathrm{smp}}]$, then it is reasonable to deduce that $p^{i}_{\mathrm{DS+smp}}\approx p^{j}_{\mathrm{smp}}$ and thus also that $p^{i}_{\mathrm{DS+smp}}\approx p_{\mathbf{k^{\prime}}}$. If however $\mu[X_{\mathrm{DS+smp}}]\gtrsim\mu[X_{\mathrm{smp}}]$, then the original guess of $\hat{N}_{\mathrm{s}}$ is too low, in which case, return to Step 2 and increase $\hat{N}_{\mathrm{s}}$ (i.e., here we are using the sampler noise as a baseline/reference value).
@@
    To show that the original guess of a ‘relatively large’ $\hat{N}_{\mathrm{s}}$ was valid, choose a representative posterior $p^{i}_{\mathrm{DS+smp}}\equiv q$ to be the target distribution. Produce a set of downsampled posteriors at different, nearby $N_{\mathrm{s}}$, say, $\hat{N}_{\mathrm{s}}/2$. Then, if $\mu[Y^{q}_{\mathrm{DS+smp}}(\hat{N}_{\mathrm{s}}/2)]\approx\mu[Y^{q}_{\mathrm{DS+smp}}(\hat{N}_{\mathrm{s}})]$ and $\sigma[Y^{q}_{\mathrm{DS+smp}}(\hat{N}_{\mathrm{s}}/2)]\approx\sigma[Y^{q}_{\mathrm{DS+smp}}(\hat{N}_{\mathrm{s}})]$, the downsampled posteriors are independent of $N_{\mathrm{s}}$, otherwise, return to Step 2 and increase $\hat{N}_{\mathrm{s}}$.
@@
    Again take a representative posterior $p^{i}_{\mathrm{DS+smp}}(\hat{N}_{\mathrm{s}})\equiv q$ as the target distribution. If we show that $\mu[Y^{q}_{\mathrm{DS+smp}}]$ increases (or remains constant) as $N_{\mathrm{s}}$ decreases, the original assumption of the behaviour in Step 1 is confirmed. This behaviour is indeed borne out as we shall see over the course of the experiments detailed in the following sections. Since the conditions in Step 1 & 2 are now affirmed, and the downsampled posteriors become independent of $N_{\mathrm{s}}$, i.e., converge upon the true, original posterior $p_{\mathrm{true}}$, then $p_{\mathbf{k^{\prime}}}\approx p_{\mathrm{true}}$ and thus $p^{i}_{\mathrm{DS+smp}}(\hat{N}_{\mathrm{s}})\approx p_{\mathrm{true}}$, and we can continue to use $p^{i}_{\mathrm{DS+smp}}(\hat{N}_{\mathrm{s}})\equiv q$ as the target distribution to assess the optimal downsampling rate.
@@ 4.3.3 Accuracy threshold
Since we are simulating a noisy detector, and real-world detector data will contain a non-zero contribution of noise, the variation of posteriors induced by non-zero noise realisations is a reasonable source from which one may define a threshold between ‘acceptable’ and ‘unacceptable’ variation of the posterior. However, different noise realisations tend to shift the maximum likelihood value (by around one standard deviation), and cause slight variations in the higher moments of the distribution. Whilst the latter (the change in morphology of the posterior) is the variation of interest (since zero noise realisation with downsampling affects the posterior in a similar way), the former is the largest contributor to the total divergence of posteriors. We therefore zero the mean of the posteriors defined with non-zero noise realisations, before computing the statistics of the set of distributions.
@@
The downsampling posterior noise limit shall then be found as follows. Choose some fixed sample selection vector $\mathbf{k^{\prime}}$ with $N_{\mathrm{s}}$ 1’s. Produce a set $\{p^{i}_{\mathbf{k^{\prime}},\mathrm{phys}}\}$ of $K$ posterior estimates defined by adding different random noise realisations $\mathbf{n}$ to the (prewhitened, normalised) residuals, where
$\mathbf{n}=\mathcal{N}\left(0,\frac{1}{2\sqrt{\Delta t}}\right)\,,$ (39)
and where $\mathcal{N}\left(\mu,\sigma\right)$ represents a normal distribution with mean $\mu$ and standard deviation $\sigma$, and $\Delta t$ is the sampling interval. Now define a new, mean zeroed set $\{\hat{p}^{i}_{\mathbf{k^{\prime}},\mathrm{phys}}(\theta)\}=\{p^{i}_{\mathbf{k^{\prime}},\mathrm{phys}}(\theta-\bar{\theta}^{i}_{\mathbf{k^{\prime}},\mathrm{phys}})\}$, where $\bar{\theta}^{i}_{\mathbf{k^{\prime}},\mathrm{phys}}$ is the mean of $p^{i}_{\mathbf{k^{\prime}},\mathrm{phys}}$.
@@
Since a vanishing detector noise realisation, $\mathbf{n}=\bm{0}$, yields the expected posterior distribution over noise realisations [23, 24], and it may also be taken as a representative of the set of downsampled posteriors, the zero detector noise posterior is the ideal target distribution for comparing downsampling noise to physical noise. It must also be mean-zeroed, so the target distribution is given by $\hat{q}(\theta)=p_{\mathbf{k^{\prime}}}(\theta-\bar{\theta}^{i}_{\mathbf{k^{\prime}}})$, where $\bar{\theta}_{\mathbf{k^{\prime}}}$ is the mean of the vanishing physical noise posterior, $p_{\mathbf{k^{\prime}}}$. Then with this $\hat{q}$, and the set of $K$ mean zeroed, physical noise posteriors $\{\hat{p}^{i}_{\mathbf{k^{\prime}},\mathrm{phys}}\}$, we form the set of $K$ divergences $\hat{Y}^{\hat{q}}_{\mathrm{phys}}$ using equation (38). Then this average posterior ‘morphology divergence’ induced by different noise realisations, $\mu[\hat{Y}^{\hat{q}}_{\mathrm{phys}}]$, can be used to set an acceptable level of divergence induced by downsampling.
We will now demand the following acceptability criteria for downsampled posteriors: either
$\mu[Y^{q}_{\mathrm{DS+smp}}(N_{\mathrm{s}})]$ $\lesssim\mu[\hat{Y}^{\hat{q}}_{\mathrm{phys}}(\hat{N}_{\mathrm{s}})],\;\;\;\;\;\;\forall\;\;N_{\mathrm{s}}>N_{\mathrm{s}}^{\mathrm{opt}},\;\;\;\;\;\;$ (40a)
$\mu[Y^{q}_{\mathrm{DS+smp}}(N_{\mathrm{s}})]$ $\approx\mu[Y^{q}_{\mathrm{DS+smp}}(N^{\prime}_{\mathrm{s}})],\,\forall\;\;N_{\mathrm{s}},N^{\prime}_{\mathrm{s}}>N_{\mathrm{s}}^{\mathrm{opt}}\,,$ (40b)
or,
$\mu[Y^{q}_{\mathrm{DS+smp}}(\hat{N}_{\mathrm{s}})]$ $\approx\mu[\hat{Y}^{\hat{q}}_{\mathrm{phys}}(\hat{N}_{\mathrm{s}})]\,,$ (41a)
$\mu[Y^{q}_{\mathrm{DS+smp}}(N_{\mathrm{s}})]$ $\ll\mathcal{H}_{q}\,,\;\;\;\;\;\forall\;\;N_{\mathrm{s}}>N_{\mathrm{s}}^{\mathrm{opt}},$ (41b)
@@
for some $N_{\mathrm{s}}^{\mathrm{opt}}$, where $\mathcal{H}_{q}$ is the entropy of target distribution $q$. Assuming all the sets of divergences are approximately normally distributed, equation (40a) ensures that the posterior one obtains by downsampling has roughly the same divergence from the truth as the average divergence from the truth induced by random physical noise realisations. Then (40b) strongly indicates convergence of the posterior (that the posterior structure is independent of the number of samples used within the admissible range). The alternative criteria introduced in (4.3.3), is required when downsampling still affects the posterior morphology as much as physical noise even at the valid $\hat{N}_{\mathrm{s}}$ (this can occur when physical noise may change the location, but not significantly the morphology of the posterior). When the average divergences from the target (as functions of $N_{\mathrm{s}}$) are still very small in this case, the values used in (4.3.3) may differ by orders of magnitude, whereas what is significant is the fact that the divergence is small. The minimum $N_{\mathrm{s}}^{\mathrm{opt}}$ for which the above holds will thus be considered the optimal downsampling rate.
@@ 4.3.4 Summary
In this section (Section 5.3), we have introduced conditions, which, together, will constitute our criteria for accepting that a posterior has converged to the truth at a certain $N_{\mathrm{s}}$, to within an accuracy approximately given by the morphology variation induced by physical noise.
@@
    Use the bootstrapping procedure to obtain the target distribution $q\equiv p^{i}_{\mathrm{DS+smp}}(\hat{N}_{\mathrm{s}})\approx p_{\mathrm{true}}$, accurate to within approximately the level of posterior variation induced by the sampler. That is, choose some $\hat{N}_{\mathrm{s}}$ such that
    • $\mu[X_{\mathrm{DS+smp}}(\hat{N}_{\mathrm{s}})]\approx\mu[X_{\mathrm{smp}}(\hat{N}_{\mathrm{s}})]$ ,
    • $\sigma[X_{\mathrm{DS+smp}}(\hat{N}_{\mathrm{s}})]\approx\sigma[X_{\mathrm{smp}}(\hat{N}_{\mathrm{s}})]$ ,
    • $\mu[Y^{q}_{\mathrm{DS+smp}}(\hat{N}^{\prime}_{\mathrm{s}})]\approx\mu[Y^{q}_{\mathrm{DS+smp}}(\hat{N}_{\mathrm{s}})]$ ,
    • $\sigma[Y^{q}_{\mathrm{DS+smp}}(\hat{N}^{\prime}_{\mathrm{s}})]\approx\sigma[Y^{q}_{\mathrm{DS+smp}}(\hat{N}_{\mathrm{s}})]$,
    where $\hat{N}^{\prime}_{\mathrm{s}}\lesssim\hat{N}_{\mathrm{s}}/2$.
@@
    Then find the smallest ${N_{\mathrm{s}}}^{\mathrm{opt}}$ for which either (4.3.3) or (4.3.3) is satisfied for all $N_{\mathrm{s}}>{N_{\mathrm{s}}}^{\mathrm{opt}}$; this shall be considered the optimal downsampling rate.
@@ 4.4.1 Structure of analysis
We would like to study different sorts of posteriors that cover a range of downsampling parameters and signal types of interest, that are connected to envisaged future investigations. These include different downsampling schemes (random/hybrid/cluster), the downsampling method (single factor/FIM preservation), original number of samples in signal $N_{\mathrm{f}}$, fraction of frequency covered by the signal $\Delta f/f_{\mathrm{max}}$, where $\Delta f=f_{\mathrm{max}}-f_{\mathrm{min}}$, and of course the reduced number of samples $N_{\mathrm{s}}$. We define the space of injection signal parameters by
@@
That is, $N_{\mathrm{s}}(n)=\lfloor 2^{n/2}+\tfrac{1}{2}\rfloor$ samples, where $8\leq n\leq 26$ and where $\lfloor x+\tfrac{1}{2}\rfloor$ rounds $x$ to its nearest integer. We will fix $f_{\mathrm{max}}=0.1\,\mathrm{Hz}$ for all systems being tested. These variables are used because of their stronger connection with the downsampling procedure itself, rather than simply the morphology of the posterior that changes with changing BHB parameters.
As we need to quantify the posterior noise for each combination of system parameters, we require a number of posteriors for each combination. We will produce 21 posteriors of each combination, giving 21 divergences of the sets $X_{\mathrm{est}}$, and (${21\choose 2}=210$ divergences of the sets $Y_{\mathrm{est}}$. Thus for each downsampling scheme and each method, this equates to $3\times 3\times 22\times 21=4158$ posterior distributions. Since there are also $3\times 2=6$ combinations of scheme and method, this would yield $6\times 4158=24948$ posteriors, so to reduce the workload, we will split the analysis into two parts. The first part will use one injection signal to determine the optimal downsampling scheme and method, and the second part will use that scheme and method to determine the optimal downsampling rate over the injection signal parameter space, which will conclude our investigation.

### v2

@@ 4.3.2 Bootstrapping
    Monotonicity assumption. We assume that, for fixed system and sampler settings, the variability induced purely by downsampling is a non-increasing function of $N_{\mathrm{s}}$. Hence there should exist a minimal value $N_{\mathrm{s}}^{\min}$ such that, for all $N_{\mathrm{s}}\geq N_{\mathrm{s}}^{\min}$, the downsampling noise is acceptable in the sense that it does not exceed the pre-defined accuracy threshold $\tau$.
@@
    Candidate choice $\hat{N}_{\mathrm{s}}$ and generation of posteriors. Choose a candidate value $\hat{N}_{\mathrm{s}}$ that is expected to be “sufficiently large”. For this $\hat{N}_{\mathrm{s}}$, construct a representative target posterior
    $q(\theta)\equiv p_{\mathbf{k^{\prime}}}(\theta)$
    corresponding to some fixed sample-selection vector $\mathbf{k^{\prime}}$ with $\hat{N}_{\mathrm{s}}$ ones (for example using a zero-noise realisation, as described in Sec. 4.3.3). Then generate a set of $K$ downsampled posteriors $\{p^{i}_{\mathrm{DS+smp}}(\theta;\hat{N}_{\mathrm{s}})\}$ using different random realisations of $\mathbf{k}$ with $\hat{N}_{\mathrm{s}}$ ones.
@@
    Acceptance test for $\hat{N}_{\mathrm{s}}$. For each of the $K$ realisations, compute the divergence from the target,
    $Y^{q}_{\mathrm{DS+smp}}(\hat{N}_{\mathrm{s}})\equiv\Big\{\overline{D}_{\mathrm{KL}}\!\left(q\,\Big\|\,p^{i}_{\mathrm{DS+smp}}(\hat{N}_{\mathrm{s}})\right)\Big\}_{i=1}^{K}.$
    and form the empirical mean $\mu\big[Y^{q}_{\mathrm{DS+smp}}(\hat{N}_{\mathrm{s}})\big]$. We then compare this mean divergence with the accuracy threshold $\tau$:
    $\mu\big[Y^{q}_{\mathrm{DS+smp}}(\hat{N}_{\mathrm{s}})\big]\;\lesssim\;\tau\,.$
    If this condition is satisfied, then $\hat{N}_{\mathrm{s}}$ is deemed admissible. If not, $\hat{N}_{\mathrm{s}}$ is too small, and we increase $\hat{N}_{\mathrm{s}}$ and repeat Steps 2–3. The smallest admissible value is denoted $N_{\mathrm{s}}^{\min}$.
@@
    Posterior stability near $N_{\mathrm{s}}^{\min}$. To verify that $N_{\mathrm{s}}^{\min}$ lies in a regime where the posterior is stable, retain a representative posterior
    $q(\theta)\equiv p^{i}_{\mathrm{DS+smp}}(\theta;N_{\mathrm{s}}^{\min})$
    as the target distribution. Generate additional downsampled posteriors for nearby values $N_{\mathrm{s}}>N_{\mathrm{s}}^{\min}$ (e.g. $N_{\mathrm{s}}=2N_{\mathrm{s}}^{\min}$), compute the corresponding divergences $Y^{q}_{\mathrm{DS+smp}}(N_{\mathrm{s}})$, and check that
    $\mu\big[Y^{q}_{\mathrm{DS+smp}}(N_{\mathrm{s}})\big]$ $\approx\mu\big[Y^{q}_{\mathrm{DS+smp}}(N_{\mathrm{s}}^{\min})\big],$
    $\text{for all }N_{\mathrm{s}}>N_{\mathrm{s}}^{\min}.$
    This confirms that once $N_{\mathrm{s}}$ exceeds $N_{\mathrm{s}}^{\min}$, the posterior geometry is effectively independent of the precise choice of $N_{\mathrm{s}}$ within the admissible range.
@@
    Consistency with monotonicity and definition of $N_{\mathrm{s}}^{\mathrm{opt}}$. Finally, examine $\mu\big[Y^{q}_{\mathrm{DS+smp}}(N_{\mathrm{s}})\big]$ as a function of $N_{\mathrm{s}}$ below and above $N_{\mathrm{s}}^{\min}$. If the mean divergence increases (or remains constant) as $N_{\mathrm{s}}$ decreases below $N_{\mathrm{s}}^{\min}$, the qualitative behaviour assumed in Step 1 is confirmed. One may then choose an “optimal” downsampling rate $N_{\mathrm{s}}^{\mathrm{opt}}\geq N_{\mathrm{s}}^{\min}$, for example the smallest value at which the criteria in Eqs. (29)–(30) (Sec. 4.3.3) are satisfied.
In summary, this bootstrapping procedure uses the monotonic dependence of downsampling noise on $N_{\mathrm{s}}$ together with an externally defined accuracy threshold $\tau$ (based, in our case, on mean-zeroed physical noise realisations) to identify a minimal and, subsequently, an optimal downsampling rate.
@@ 4.3.3 Accuracy threshold
As discussed in Sec. 4.1, different realisations of detector noise induce variability in the posterior that we regard as irreducible in realistic data analyses. We are interested only in changes to the geometry of the posterior, not in noise-induced shifts of its mean, so we work with mean-zeroed posteriors when quantifying these effects. Using this setup, we now define an accuracy threshold $\tau$ based on the typical geometry variation induced by physical noise.
The downsampling posterior-noise limit is then found as follows.
  1. 1.
    Choose some sample-selection vector $\mathbf{k^{\prime}}$ with $N_{\mathrm{s}}$ ones.
  2. 2.
@@
    Produce a set $\{p^{i}_{\mathbf{k^{\prime}},\mathrm{phys}}\}$ of $K$ posterior estimates defined by adding different random noise realisations to the (prewhitened, normalised) residuals.
  3. 3.
    Define a new, mean-zeroed set
    $\left\{\hat{p}^{i}_{\mathbf{k^{\prime}},\mathrm{phys}}(\bm{\theta})\right\}=\left\{p^{i}_{\mathbf{k^{\prime}},\mathrm{phys}}(\bm{\theta}-\bar{\bm{\theta}}^{i}_{\mathbf{k^{\prime}},\mathrm{phys}})\right\},$
    where $\bar{\bm{\theta}}^{i}_{\mathbf{k^{\prime}},\mathrm{phys}}$ is the vector of posterior means of $p^{i}_{\mathbf{k^{\prime}},\mathrm{phys}}$.
@@
Since a vanishing detector-noise realisation, $\mathbf{n}=\bm{0}$, yields the expected posterior distribution over noise realisations [GPULISA, AvgNoise], and may also be taken as representative of the set of downsampled posteriors, the zero-noise posterior is the ideal target distribution for comparing downsampling noise to physical noise. It must also be mean-zeroed, so we define the target distribution
$\hat{q}(\bm{\theta})=p_{\mathbf{k^{\prime}}}(\bm{\theta}-\bar{\bm{\theta}}_{\mathbf{k^{\prime}}}),$
where $\bar{\bm{\theta}}_{\mathbf{k^{\prime}}}$ is the mean of the zero-noise posterior $p_{\mathbf{k^{\prime}}}$. With this $\hat{q}$, and the set of $K$ mean-zeroed physical-noise posteriors $\{\hat{p}^{i}_{\mathbf{k^{\prime}},\mathrm{phys}}\}$, we form the set of $K$ divergences $\hat{Y}^{\hat{q}}_{\mathrm{phys}}$ using Eq. (27). The resulting average posterior “geometry divergence” induced by different noise realisations, $\mu[\hat{Y}^{\hat{q}}_{\mathrm{phys}}$], is then used to set an acceptance threshold for divergence induced by downsampling.
We define the accuracy threshold
$\tau\;\equiv\;\mu\big[\hat{Y}^{\hat{q}}_{\mathrm{phys}}\big$]\,. (28)
This $\tau$ is then used in Sec. 4.3.2 as the reference level against which downsampling-induced variability is judged.
We now demand the following acceptability criteria for downsampled posteriors: either
$\mu\big[Y^{q}_{\mathrm{DS+smp}}(N_{\mathrm{s}})\big]$ $\lesssim\tau,$ (29a)
$\mu\big[Y^{q}_{\mathrm{DS+smp}}(N_{\mathrm{s}})\big]$ $\approx\mu\big[Y^{q}_{\mathrm{DS+smp}}(N^{\prime}_{\mathrm{s}})\big]\,$, (29b)
for all $N_{\mathrm{s}},N^{\prime}_{\mathrm{s}}>N_{\mathrm{s}}^{\mathrm{opt}}$, or,
$\mu\big[Y^{q}_{\mathrm{DS+smp}}(\hat{N}_{\mathrm{s}})\big]$ $\approx\tau\,$, (30a)
$\mu\big[Y^{q}_{\mathrm{DS+smp}}(N_{\mathrm{s}})\big]$ $\ll\mathcal{H}_{q}\,$, (30b)
for all $N_{\mathrm{s}}>N_{\mathrm{s}}^{\mathrm{opt}}$, for some $N_{\mathrm{s}}^{\mathrm{opt}}$, where $\mathcal{H}_{q}$ is the entropy of the target distribution $q$.
Even though the empirical distributions of divergences are non-negative and typically skewed, $\mu[X]$ remains a convenient summary statistic. Equation (29a) ensures that the posterior obtained by downsampling has roughly the same divergence from the target as the average divergence induced by random physical noise realisations. Condition (29b) then strongly indicates convergence of the posterior, in the sense that the posterior structure is insensitive to the number of samples used within the admissible range.
@@
The alternative criteria in (30) are required when downsampling still affects the posterior geometry as much as physical noise even at the candidate $\hat{N}_{\mathrm{s}}$ (this can occur when physical noise mainly shifts the location but does not significantly change the geometry of the posterior). In such cases the average divergences from the target as functions of $N_{\mathrm{s}}$ may be very small, and the absolute values in (29) can differ by orders of magnitude; what matters is that the divergence remains small compared to the intrinsic information scale $\mathcal{H}_{q}$. The minimum $N_{\mathrm{s}}^{\mathrm{opt}}$ for which the above holds is then taken to be the optimal downsampling rate.
@@ 4.3.4 Summary
In this section we have introduced a set of conditions which, taken together, constitute our acceptance criteria for declaring that a downsampled posterior has converged to the true posterior at a given $N_{\mathrm{s}}$, up to an accuracy comparable to the geometry variation induced by physical detector noise.
@@
    Instantiate the accuracy threshold $\tau$ using physical-noise realisations as described in Sec. 4.3.3, i.e. set $\tau\equiv\mu[\hat{Y}^{\hat{q}}_{\mathrm{phys}}$] for the corresponding mean-zeroed physical-noise divergences.
@@
    Use the bootstrapping procedure of Sec. 4.3.2 with this accuracy threshold $\tau$ to identify a minimal admissible downsampling rate $N_{\mathrm{s}}^{\min}$ and a corresponding target distribution $q\approx p_{\mathrm{true}}$, chosen as a representative downsampled posterior at $N_{\mathrm{s}}^{\min}$, such that
    • the mean divergence from $q$ at $N_{\mathrm{s}}^{\min}$ satisfies $\mu\big[Y^{q}_{\mathrm{DS+smp}}(N_{\mathrm{s}}^{\min})\big]\lesssim\tau$, and
    • for all $N_{\mathrm{s}}>N_{\mathrm{s}}^{\min}$, the mean divergence is stable, $\mu\big[Y^{q}_{\mathrm{DS+smp}}(N_{\mathrm{s}})\big]\approx\mu\big[Y^{q}_{\mathrm{DS+smp}}(N_{\mathrm{s}}^{\min})\big]$.
  3. 3.
    Finally, within the admissible regime $N_{\mathrm{s}}\geq N_{\mathrm{s}}^{\min}$, find the smallest $N_{\mathrm{s}}^{\mathrm{opt}}$ for which either the criteria in Eqs. (29) or those in Eqs. (30) are satisfied for all $N_{\mathrm{s}}>N_{\mathrm{s}}^{\mathrm{opt}}$. This $N_{\mathrm{s}}^{\mathrm{opt}}$ is then taken to be the optimal downsampling rate.
@@ 4.4.1 Structure of analysis
We study posteriors across a range of downsampling parameters and signal types, chosen to be representative of envisaged future LISA analyses. These include different downsampling schemes (random/hybrid/cluster), the downsampling method (single factor/FIM preservation), original number of samples in signal $N_{\mathrm{f}}$, fraction of frequency covered by the signal $\Delta f/f_{\mathrm{max}}$, where $\Delta f=f_{\mathrm{max}}-f_{\mathrm{min}}$, and of course the reduced number of samples $N_{\mathrm{s}}$. We define the space of injection signal parameters by
@@
That is, $N_{\mathrm{s}}(n)=\lfloor 2^{n/2}+\tfrac{1}{2}\rfloor$ samples, where $8\leq n\leq 26$ and where $\lfloor x+\tfrac{1}{2}\rfloor$ rounds $x$ to its nearest integer. We will fix $f_{\mathrm{max}}=0.1\,\mathrm{Hz}$ for all systems being tested. This parametrisation of the fiducial systems is chosen to probe how the downsampling procedure behaves as a function of generic signal properties (e.g. frequency span in the Fourier domain and susceptibility to aliasing or degradation under downsampling).
As we need to quantify the posterior noise for each combination of system parameters, we require a number of posteriors for each combination. We produce 21 posteriors per combination, giving ${21\choose 2}=210$ divergences in the sets $X_{\mathrm{est}}$, and 21 divergences in the sets $Y^{q}_{\mathrm{est}}$. Thus, for each downsampling scheme and each method, this equates to $3\times 3\times 22\times 21=4158$ posterior distributions. Since there are also $3\times 2=6$ combinations of scheme and method, this would yield $6\times 4158=24948$ posteriors in total. To reduce the workload, we therefore split the analysis into two parts. The first part uses a single injection signal to determine the optimal downsampling scheme and method; the second part uses that scheme and method to determine the optimal downsampling rate over the injection-signal parameter space.